simplify the expression using the quotients to power rule. If possible, evaluate exponential expressions.
-3a to the third power/b to the second power and both of the to the power of 4

Answer by dabanfield(803)

simplify the expression using the quotients to power rule. If possible, evaluate exponential expressions.
-3a to the third power/b to the second power and both of the to the power of 4
((-3a)^3/(b^2))^4 =
(-3a)^3)^4/(b^2)^4 =
(-3a)^12/b^8 =
((-3)^12*a^12)/b^8